# RELATE: Relation Extraction in Biomedical Abstracts with LLMs and Ontology Constraints

### v1

@@ Abstract
Biomedical knowledge graphs (KGs) are vital for drug discovery and clinical decision support but remain incomplete. Large language models (LLMs) excel at extracting biomedical relations, yet their outputs lack standardization and alignment with ontologies, limiting KG integration. We introduce RELATE, a three-stage pipeline that maps LLM-extracted relations to standardized ontology predicates using ChemProt and the Biolink Model. The pipeline includes: (1) ontology preprocessing with predicate embeddings, (2) similarity-based retrieval enhanced with SapBERT, and (3) LLM-based reranking with explicit negation handling. This approach transforms relation extraction from free-text outputs to structured, ontology-constrained representations. On the ChemProt benchmark, RELATE achieves 52% exact match and 94% accuracy@10, and in 2,400 HEAL Project abstracts, it effectively rejects irrelevant associations (0.4%) and identifies negated assertions. RELATE captures nuanced biomedical relationships while ensuring quality for KG augmentation. By combining vector search with contextual LLM reasoning, RELATE provides a scalable, semantically accurate framework for converting unstructured biomedical literature into standardized KGs.
@@
Relation Extraction, Large Language Models, Ontology, Biolink Protocol, Biomedical Knowledge Graphs
@@
Data used in this project are publicly available ChemProt dataset [^1] https://huggingface.co/datasets/bigbio/chemprot and PubMed abstracts available through the NIH Helping to End Addiction Long-term (HEAL) Initiative [^2] https://heal.nih.gov/. The code is available in a public Git repository [^3] https://github.com/RENCI-NER/pred-mapping/tree/multi-ontology.
@@ 1 Introduction
Knowledge graphs (KGs) serve as structured repositories of facts, capturing entities and their relationships to support applications such as semantic search, recommendation systems, and drug repurposing (Olasunkanmi et al., 2024; Bizon et al., 2019). However, real-world KGs are often incomplete, with missing facts that limit their utility.
@@
Recent advancements in large language models (LLMs) have demonstrated remarkable capabilities in biomedical text understanding (Jahan et al., 2024). Nevertheless, current LLM-based RE methods often lack integration with domain ontologies, leading to high false-positive rates and inconsistent relationship representations. Standardized ontology frameworks such as the Biolink Model (Unni et al., 2022) define well-structured predicates for biomedical relationships (Bizon et al., 2019; Joachimiak et al., 2021; Reese et al., 2021; all, 2022). Yet, mapping free-text relations to ontology predicates remains a difficult task due to the semantic gap between natural language expressions and formal ontological representations.
@@
    Ontology-constrained relation standardization: RELATE introduces the first systematic framework for mapping LLM-extracted biomedical relations to established ontologies (ChemProt and Biolink), moving beyond relation classification toward ontology-grounded standardization. This design ensures semantic interoperability with biomedical KGs. While ChemProt has been used for supervised relation classification (Peng et al., 2019; Warikoo et al., 2021), and Biolink for knowledge graph schema definition (Joachimiak et al., 2021; Reese et al., 2021), our work represents the first application of these ontologies as standardization targets for free-text relation extraction, bridging the gap between unstructured text mining and structured knowledge representation.
@@
    Comprehensive evaluation: We conduct dual-ontology assessment on ChemProt (Peng et al., 2019) benchmark and 2,400 real-world HEAL Project abstracts. Integration with an established biomedical KG (ROBOKOP) is also planned to further demonstrate its translational impact.
@@
Figure 1: The three-stage RELATE pipeline: (1) ontology preprocessing, (2) similarity-based retrieval, and (3) contextual refinement. Ontology preprocessing generates embeddings for predicates and their negated variants. It updates only if the ontology schema or embedding models change. Given an input triple with abstract context, RELATE performs similarity-based retrieval by embedding the relation and retrieving top-$k$ ontology candidates. In the final stage, contextual refinement reranks these candidates with an LLM using the full abstract context, producing the most semantically appropriate ontology predicate.
@@ 2 Related Work
No existing framework systematically bridges free-text biomedical RE with standardized ontological predicates. While zero-shot LLMs (Li et al., 2025) can outperform fine-tuned models on smaller datasets, they struggle when confronted with complex, domain-specific tasks that demand specialized knowledge (Chen et al., 2024a; Jahan et al., 2024). Further research has explored adaptive instruction-rich prompting (Zhou et al., 2024; Tao et al., 2024) and soft prompt-based learning approaches that allow models to automatically optimize prompts rather than relying on hand-crafted instructions (Peng et al., 2024). These methods show promise but have not achieved systematic biomedical ontology integration.
@@
Fine-tuning approaches have demonstrated strong performance on domain-specific benchmarks. Techniques such as adaptive document mapping, ensemble learning with attention mechanisms, and specialized training procedures (Shang et al., 2025; Zhong et al., 2024; Orlova and Orlov, 2025) improve classification accuracy. However, they require extensive retraining and do not address the core challenge of standardizing predicates extracted from free-text relations.
@@
Recent work in LLM-based relation extraction has explored various validation and standardization approaches. RelCheck (Ourekouch et al., 2025) addresses low-confidence predictions from pretrained models using ontology-guided LLM validation, though it focuses on confidence rather than standardization. SPIREX (Cavalleri et al., 2024) applies schema-constrained prompts with graph machine learning validation but restricts itself to RNA-specific domains. Graph-augmented approaches such as RAG-enhanced LLMs for automated ontology extension (Georgakopoulos et al., 2025); structure-oriented RAG for KG-LLM co-learning (Yang et al., 2024); and integrated GNNs with LLM-generated support documents (Dong et al., 2024), and LLM-driven ontology enrichment pipelines (Kollapally et al., 2025).
However, these approaches primarily target domain-specific extraction, confidence validation, or knowledge system extension rather than systematic standardization of extracted relationships to established ontological frameworks. Multi-channel neural architectures combine textual and graph structures but focus primarily on structural features rather than systematic predicate mapping. Caufield et al. (2024) applied structured prompt interrogation for ontology-grounded annotations from scientific publications, yet their focus remained on entity linking rather than comprehensive relation standardization. Even enhancements to systems such as SemRep (Ming et al., 2024) that leverage contrastive learning operate within existing predicate frameworks rather than enabling mapping to broader ontological schemas.
@@ 3.1 Stage 1: Ontology Preprocessing
The first stage of our pipeline (Alg. 1) extracts predicates with positive and negative descriptors to encode domain knowledge, and transforms them into a searchable embedding database suitable for relation mapping.
@@
We begin by collecting all predicate definitions from biomedical ontologies such as the Biolink Model toolkit (Unni et al., 2022) or ChemProt (Krallinger et al., 2017), thereby constructing a base vocabulary of biomedical relationships. We will denote the set of all ontology predicates as $\mathcal{R}$. Each predicate $r\in\mathcal{R}$ represents a distinct biological or medical relationship (e.g., affects, inhibits) that can exist between entities in the KG.
@@
The descriptor dataset $\mathcal{D}_{r}^{+}$ is constructed by gathering existing textual descriptions for each predicate $r$. For example, one descriptor for the predicate affects may be “describes an entity that has an effect on the state or quality of another existing entity”. These descriptions are sourced from multiple authoritative biomedical sources. In the Biolink example, this includes official documentation and standardized vocabularies, while in the ChemProt case, online relation category descriptions suffice. The distributions of the predicate descriptors are presented in Table LABEL:tab:top_predicates_distribution_chemprot and Table LABEL:tab:top_predicates_distribution_heal. These aggregated descriptions form the positive dataset $\mathcal{D}_{r}^{+}$ and capture the semantic diversity with which biomedical relationships appear in scientific literature.
@@
To complement the positive dataset, we introduce a novel strategy for generating negative descriptors. For each descriptor in $D_{r}^{+}$, an LLM produces natural, semantically coherent negations using a structured prompt (Figure LABEL:fig:negation_prompt). For example, the earlier descriptor becomes “describes an entity that does not have an effect on the state or quality of another existing entity.” These negated descriptors are stored in a separate dictionary $D_{r}^{-}$ to serve as contrastive examples that help the embedding model distinguish valid from invalid predicate mappings. $D_{r}^{-}$ is labeled by $r$ + “_NEG” (i.e. affects $\to$ affects_NEG) to maintain clear separation from positive instances.
@@
Finally, every descriptor in $D^{+}_{r}$ and $D^{-}_{r}$ is transformed into $d$-dimensional vector representations using the embedding model $f_{\theta}$. Instead of aggregating these descriptor embeddings to a single predicate embedding, we store all descriptor embeddings, $\mathcal{V}_{r}$, in a searchable database for the next stage. To address limitations of general-purpose embedding models (e.g., nomic-text-embed or bge-m3) in biomedical domains, we introduce an (optional) hybrid retrieval enhancement based on SapBERT (Liu et al., 2021), a model specialized in biomedical entity linking. In this setup (Figure LABEL:fig:workflowplus), SapBERT was fine-tuned on $D^{+}_{r}$ and $D^{-}_{r}$ to generate a second set of embeddings $\mathcal{V}^{\prime}_{r}$. While SapBERT has demonstrated effectiveness in biomedical entity alignment (Liu et al., 2021), this work presents its first application to relation extraction through predicate embedding enhancement. Our hybrid approach leverages SapBERT’s biomedical domain knowledge to improve relation semantic similarity beyond general-purpose embedding models.
@@
Figure 2: Negation generation prompt used in Stage 3.1 for creating negative descriptors from positive descriptors.
@@ 3.2 Stage 2: Similarity-Based Retrieval
For every predicate $r\in\mathcal{R}$, Stage 1 has produced positive and negative descriptor dictionaries $\mathcal{D}_{r}^{+}$, $\mathcal{D}_{r}^{-}$ and corresponding descriptor embedding vectors $\mathcal{V}_{r}$. We now leverage these data to retrieve candidate ontology predicates for incoming extracted relations.
The extracted relations come in a quadruple $(s,o,T,a)$: $s$ subject entity, $o$ object entity, $T$ the free-form relation text, and $a$ the abstract context from which the relation was identified. For example, $T$ could be “increases level of,” which does not exist in Biolink ontology in this format. First, free-form relation text $T$ is embedded to query vector $q=f_{\theta}(T)$. We use the same model $f_{\theta}$ from Stage 1, ensuring semantic consistency between query relations and stored ontology predicates.
@@
By narrowing the search space from hundreds of ontology predicates to a small candidate set, this stage provides an efficient yet semantically informed mechanism for predicate selection. At this point, however, the search operates without the benefit of contextual information from the original biomedical abstract, focusing only on lexical and semantic similarity.
@@ 3.3 Stage 3: Contextual Refinement
- If no match exists, return ‘"mapped_predicate": "none"‘.
@@
{{"mapped_predicate": "one of the predicate keys or ’none’", "negated": "True" or "False"}}
Figure 3: Contextual reranking prompt used in Stage 3 for LLM-based predicate selection with explicit negation handling.
The final stage addresses the key limitation of similarity-based retrieval: its reliance on context-free embeddings. While Stage 2 efficiently narrows the predicate space to a small candidate set, it does not incorporate the biomedical context in which the relation was expressed. Stage 3 introduces contextual refinement, where an LLM evaluates and re-ranks the top-$k$ candidates against the full relation context.
@@
The LLM then performs re-ranking of the candidate set, selecting the single predicate that best aligns with the domain context. Importantly, the system includes a NONE option, allowing the model to reject all candidates if none provide a semantically valid match. This safeguard prevents the forced assignment of predicates in cases where the extracted relation is either too domain-specific, erroneous, or incompatible with the ontology.
The output of Stage 3 is a contextually validated predicate $r^{*}$ that integrates both semantic similarity and biomedical context. This refinement ensures that only relationships consistent with both the ontology and the source text are admitted into the knowledge graph, thereby improving accuracy and reducing false positives.

### v2

@@ Abstract
Biomedical knowledge graphs (KGs) are vital for drug discovery and clinical decision support but remain incomplete. Large language models (LLMs) excel at extracting biomedical relations, yet their outputs lack standardization and alignment with ontologies, limiting KG integration with free texts. We introduce RELATE, a three-stage pipeline that maps LLM-extracted relations to standardized ontology predicates, e.g., the Biolink Model. The pipeline includes: (1) ontology preprocessing with predicate embeddings, (2) similarity-based retrieval enhanced with SapBERT, and (3) LLM-based reranking with explicit negation handling. This approach performs relation extraction from free-text outputs to structured, ontology-constrained representations. On the ChemProt benchmark, RELATE achieves 52% exact match and 94% accuracy@10, and in 2,400 HEAL Project abstracts, it effectively rejects irrelevant associations (0.4%) and identifies negated assertions. RELATE captures nuanced biomedical relationships while ensuring quality for KG augmentation. By combining vector search with contextual LLM reasoning, RELATE provides a scalable, semantically accurate framework for converting unstructured biomedical literature into standardized KGs.
@@
Relation Extraction, Large Language Models, Ontology, Biolink Model, Biomedical Knowledge Graphs
@@
Data used in this project are publicly available ChemProt dataset [^1] https://huggingface.co/datasets/bigbio/chemprot and PubMed abstracts available through the NIH Helping to End Addiction Long-term (HEAL) Initiative [^2] https://heal.nih.gov/. The code is available in a public Git repository [^3] https://github.com/RENCI-NER/pred-mapping/releases/tag/v1.0.
@@ 1 Introduction
Knowledge graphs (KGs) serve as structured repositories of facts, capturing entities and their relationships to support applications such as semantic search, recommendation systems, and drug repurposing (olasunkanmi2024explainable; bizon2019robokop). However, real-world KGs are often incomplete, with missing facts that limit their utility.
@@
Recent advancements in large language models (LLMs) have demonstrated remarkable capabilities in biomedical text understanding (jahan2024comprehensive). Nevertheless, current LLM-based RE methods often lack integration with domain ontologies, leading to high false-positive rates and inconsistent relationship representations. Standardized ontology frameworks such as the Biolink Model (unni2022biolink) define well-structured predicates for biomedical relationships (bizon2019robokop; joachimiak2021kg; reese2021kg; alliance2022harmonizing). Yet, mapping free-text relations to ontology predicates remains a difficult task due to the semantic gap between natural language expressions and formal ontological representations.
@@
    Ontology-constrained relation standardization: RELATE introduces the first systematic framework for mapping LLM-extracted biomedical relations to established ontologies (ChemProt and Biolink), moving beyond relation classification toward ontology-grounded standardization. This design ensures semantic interoperability with biomedical KGs. While ChemProt has been used for supervised relation classification (peng2019transfer; warikoo2021lbert), and Biolink for KG schema definition (joachimiak2021kg; reese2021kg), our work represents the first application of these ontologies as standardization targets for free-text relation extraction, bridging the gap between unstructured text mining and structured knowledge representation.
@@
    Comprehensive evaluation: We conduct dual-ontology assessment on ChemProt (peng2019transfer) benchmark and 2,400 real-world HEAL Project abstracts. Integration with an established, large-scale biomedical KG (ROBOKOP) is also planned to further demonstrate its translational impact.
These capabilities enable downstream applications in literature-based discovery, biomedical KG construction, and multi-source knowledge integration; see Appendix A.
@@
Figure 1: The three-stage RELATE pipeline: (1) ontology preprocessing, (2) similarity-based retrieval, and (3) contextual refinement. Ontology preprocessing generates embeddings for predicates and their negated variants. It updates only if the ontology schema or embedding models change. Given an input quadruple—subject, object, relation text, and abstract context—RELATE performs similarity-based retrieval by embedding the relation and retrieving top-$k$ ontology candidates. In the final stage, contextual refinement reranks these candidates with an LLM using the full abstract context, producing the most semantically appropriate ontology predicate.
@@ 2 Related Work
No existing framework systematically bridges free-text biomedical relation extraction (RE) with standardized ontological predicates. While zero-shot LLMs (li2025grag) can outperform fine-tuned models on smaller datasets, they struggle when confronted with complex, domain-specific tasks that demand specialized knowledge (chen2024benchmarking; jahan2024comprehensive). Further research has explored adaptive instruction-rich prompting (zhou2024leap; tao2024graphical) and soft prompt-based learning approaches that allow models to automatically optimize prompts rather than relying on hand-crafted instructions (peng2024model). These methods show promise but have not achieved systematic biomedical ontology integration.
@@
Fine-tuning approaches have demonstrated strong performance on domain-specific benchmarks. Techniques such as adaptive document mapping, ensemble learning with attention mechanisms, and specialized training procedures (shang2025biomedical; zhong2024leveraging; orlova2025challenges) improve classification accuracy. However, they require extensive retraining and do not address the core challenge of standardizing predicates extracted from free-text relations. chandak2023building emphasized the importance of standardized schemas and quality control in its PrimeKG framework by integrating heterogeneous data sources. RELATE applies these principles to the RE layer, enabling automated standardization of literature-derived relations for seamless integration into KGs such as PrimeKG, with ensured schema compliance and semantic interoperability.
@@
Recent work in LLM-based RE has explored various validation approaches. RelCheck (ourekouch2025relcheck) addresses low-confidence predictions from pretrained models using ontology-guided LLM validation, though it focuses on confidence rather than standardization. SPIREX (cavalleri2024spirex) applies schema-constrained prompts with graph machine learning validation but restricts itself to RNA-specific domains. Graph-augmented approaches include RAG-enhanced LLMs for automated ontology extension (georgakopoulos2025text), structure-oriented RAG for KG-LLM co-learning (yang2024knowledge), integrated GNNs with LLM-generated support documents (dong2024graph), and LLM-driven ontology enrichment pipelines (kollapally2025ontology). However, these approaches primarily target domain-specific extraction or confidence validation rather than systematic standardization of extracted relationships to established ontological frameworks. Multi-channel neural architectures combine textual and graph structures but focus primarily on structural features rather than systematic predicate mapping. caufield2024structured used structured prompts for ontology-grounded annotations, focusing on entity linking rather than full relation standardization, while enhancements to SemRep (ming2024enhancing) remain within existing predicate frameworks.
While cao2024automatic built an end-to-end system for extracting knowledge about rare diseases, their work addresses only 6 relation types in a narrow domain. More fundamentally, no existing framework systematically bridges free-text biomedical RE with standardized ontological predicates. RELATE addresses this critical gap by mapping free-text relations to standardized ontology predicates, handling both narrow (ChemProt’s 9 predicates) and broad (Biolink’s 284 predicates) ontologies, ensuring semantic interoperability and enabling seamless integration into biomedical KGs.
@@ 3.1 Stage 1: Ontology Preprocessing
The first stage of our pipeline (Alg. 1) extracts ontology predicates with positive and negative descriptors to encode domain knowledge, and transforms them into a searchable embedding database suitable for relation mapping.
@@
We begin by collecting all predicate definitions from biomedical ontologies such as the Biolink Model toolkit (unni2022biolink) or ChemProt (DBLP:journals/biodb/LiSJSWLDMWL16), thereby constructing a base vocabulary of biomedical relationships. We will denote the set of all ontology predicates as $\mathcal{R}$. Each predicate $r\in\mathcal{R}$ represents a distinct biological or medical relationship (e.g., affects, inhibits) that can exist between entities in the KG.
@@
The descriptor dataset $\mathcal{D}_{r}^{+}$ is constructed by gathering existing textual descriptions for each predicate $r$. For example, one descriptor for the predicate affects may be “describes an entity that has an effect on the state or quality of another existing entity.” These descriptions are sourced from multiple authoritative biomedical sources. In the Biolink example, this includes official documentation and standardized vocabularies, while in the ChemProt case, online relation category descriptions suffice. The distributions of the predicate descriptors are presented in Tables LABEL:tab:top_predicates_distribution_chemprot and LABEL:tab:top_predicates_distribution_heal. These aggregated descriptions form the positive dataset $\mathcal{D}_{r}^{+}$ and capture the semantic diversity with which biomedical relationships appear in scientific literature.
@@
To complement the positive dataset, we introduce a novel strategy for generating negative descriptors. For each descriptor in $D_{r}^{+}$, an LLM produces a natural, semantically coherent negation using a structured prompt (Figure LABEL:fig:negation_prompt). For example, the earlier descriptor becomes “describes an entity that does not have an effect on the state or quality of another existing entity.” These negated descriptors are stored in a separate dictionary $D_{r}^{-}$ to serve as contrastive examples that help the embedding model distinguish valid from invalid predicate mappings. $D_{r}^{-}$ is labeled by $r$ + “_NEG” (i.e. affects $\to$ affects_NEG) to maintain clear separation from positive instances.
Finally, every descriptor in $D^{+}_{r}$ and $D^{-}_{r}$ is transformed into a $d$-dimensional vector representation using the embedding model $f_{\theta}$. Instead of aggregating these descriptor embeddings to a single predicate embedding, we store all descriptor embeddings, $\mathcal{V}_{r}$, in a searchable database for the next stage.
@@
To address limitations of general-purpose embedding models (e.g., nomic-text-embed or bge-m3) in biomedical domains, we introduce an (optional) hybrid retrieval enhancement based on SapBERT (liu2021self), a model specialized in biomedical entity linking. In this setup (Figure LABEL:fig:workflowplus), SapBERT was fine-tuned on $D^{+}_{r}$ and $D^{-}_{r}$ to generate a second set of embeddings $\mathcal{V}^{\prime}_{r}$. While SapBERT has demonstrated effectiveness in biomedical entity alignment, this work presents its first application to RE through descriptor embedding enhancement. Our hybrid approach leverages SapBERT’s biomedical domain knowledge to improve relation semantic similarity beyond general-purpose embedding models.
@@
Figure 2: Negation generation prompt used in Stage 1 (Section 3.1) for creating negative descriptors from positive descriptors.
@@ 3.2 Stage 2: Similarity-Based Retrieval
For every predicate $r\in\mathcal{R}$, Stage 1 has produced positive and negative descriptor dictionaries $\mathcal{D}_{r}^{+}$, $\mathcal{D}_{r}^{-}$ and corresponding descriptor embedding vectors $\mathcal{V}_{r}$. We now leverage these data to retrieve candidate ontology predicates for incoming relations extracted by an upstream RE system, which could be any general-purpose LLM.
The extracted relations come in a quadruple $(s,o,T,a)$: $s$ subject entity, $o$ object entity, $T$ the free-form relation text, and $a$ the abstract context from which the relation was identified. For example, in Figure LABEL:fig:workflow, $T$ is “decreases transporting activities of,” which does not exist in Biolink ontology $\mathcal{R}$ in this format.
First, free-form relation text $T$ is embedded to a query vector $q=f_{\theta}(T)$. We use the same model $f_{\theta}$ from Stage 1, ensuring semantic consistency between query relations and stored ontology predicates.
@@
By narrowing the search space from hundreds of ontology predicates to a small candidate set, this stage provides an efficient yet semantically informed mechanism for predicate selection. At this point, however, the search operates without the benefit of contextual information from the original biomedical abstract $a$, focusing only on lexical and semantic similarity.
@@ 3.3 Stage 3: Contextual Refinement
- If no match exists, return ‘"mapped_predicate": "none"’.
@@
{{"mapped_predicate": "one of the predicate keys or ‘none’", "negated": "True" or "False"}}
Figure 3: Contextual reranking prompt used in Stage 3 (Section 3.3) for LLM-based predicate selection with explicit negation handling.
The final stage addresses the key limitation of similarity-based retrieval: its reliance on context-free embeddings. While Stage 2 efficiently narrows the predicate space to a small candidate set, it does not incorporate the biomedical context in which the relation was expressed. Stage 3 introduces contextual refinement, where an LLM evaluates and reranks the top-$k$ candidates against the full relation context.
@@
The LLM then performs reranking of the candidate set, selecting the single predicate that best aligns with the domain context. Importantly, the system includes a NONE option, allowing the model to reject all candidates if none provide a semantically valid match. This safeguard prevents the forced assignment of predicates in cases where the extracted relation is either too domain-specific, erroneous, or incompatible with the ontology.
The output of Stage 3 is a contextually validated predicate $r^{*}$ that integrates both semantic similarity and biomedical context. This refinement ensures that only relationships consistent with both the ontology and the source text are admitted into the KG, thereby improving accuracy and reducing false positives.
@@ 4 Experiments
We assess RELATE on two independent datasets with different ontology models, as well as four different embedding model configurations. See Appendix D for example descriptors in both datasets.
4.1 Datasets
The ChemProt dataset (DBLP:journals/biodb/LiSJSWLDMWL16) provides annotated chemical-protein relations representing diverse pharmacological and biochemical relationship types, including regulators, modulators, agonists, antagonists, and substrate interactions. We treat the information in ChemProt as ground truth abstract-predicate pairs for evaluation purposes. Standard metrics in information retrieval are used, such as exact match accuracy, accuracy@$k$, and Mean Reciprocal Rank (MRR). Recall from Figure LABEL:fig:reranking_prompt that Stage 3 of RELATE only outputs its top choice predicate $r^{*}$ or NONE, and does not return a fully reranked or reordered list of $k$ candidates. Therefore, exact match accuracy is calculated with respect to the final output of RELATE ($r^{*}$), but accuracy@$k$ and MRR are calculated using the top-$k$ candidates after Stage 2, before LLM-based Stage 3. This allows us to analyze the impact of LLM-based refinement in isolation.
Next, we apply RELATE to 2,400 PubMed abstracts from the HEAL Project [^4] https://heal.nih.gov targeting opioid use disorder research. This real-world dataset reflects the practical scenario in which extracted relationships must be mapped to standardized Biolink predicates for integration into a large biomedical KG [^5] https://robokop.renci.org/. Unlike ChemProt, the HEAL abstracts lack ground-truth annotations, requiring qualitative assessment of RELATE’s outputs.
Importantly, this corpus captures the complexity and noisiness of real biomedical literature. In contrast to benchmark datasets, these abstracts often contain methodological statements or negative scientific findings (e.g., “drug does not affect condition,” “gene not associated with disease”). Such cases underscore the necessity of rejection and negation-handling to ensure only semantically appropriate biomedical relationships enter the KG.
4.2 LLM and Embedding Models
We experimented with MedGemma (4B), LLaMA2 (7B), LLaMA3 (8B), LLaMA3 (70B), MedGemma (27B), and OpenAI GPT-4o-mini (unknown number of parameters). The last two demonstrated greater consistency and fewer hallucinations compared to the other models. Since those two models had very similar performances, we only report results using MedGemma (27B) in this work. However, note that any other LLM could be substituted to reproduce the pipeline. MedGemma (27B) is Google’s medical language model (sellergren2025medgemma), trained exclusively on medical text and optimized for efficient inference through architectural improvements in performance and computational efficiency.
To evaluate RELATE’s pipeline contribution beyond raw LLM capability, we implemented a direct classification baseline. This baseline uses MedGemma (27B) to classify ChemProt relations by providing the same input quadruple (subject, object, relation text, abstract) along with the list of 9 ChemProt predicates. This represents a common approach in LLM-based relation extraction and provides a fair comparison since it uses the same LLM and has access to the same predicate definitions.
As with the LLMs, we also experimented with several embedding approaches. While we prioritize accessibility through ollama-based open-access models, we tested OpenAI’s text-embed-large models and others with 8,192-token context windows that provide general-purpose semantic understanding. These include nomic-embed-text (nussbaumnomic), a 768-dimensional embedding model, and bge-m3, a 1,024-dimensional multilingual embedding model for diverse textual contexts (chen2024bge). Since nomic-embed-text and bge-m3 achieved comparable performance, we report only the results obtained with the 768-dimensional nomic-embed-text in this work.
For (optional) hybrid retrieval, we experimented with adding fine-tuned SapBERT (liu2021self). We initialized the model with BiomedNLP-PubMedBERT-base-uncased-abstract -fulltext and optimized it using a contrastive learning framework. Specifically, we employed multi-similarity loss with hard negative mining, which encourages the model to better distinguish between semantically close but incorrect pairs. To keep training efficient while still capturing relevant semantic features, we limited input sequences to 25 tokens and represented each instance using the [CLS] embedding.